OpenStudy (anonymous):

equivalent logarithmic form 5^2 = x evaluate log4 64

OpenStudy (anonymous):

log(5)x=2

OpenStudy (anonymous):

log44^3=3

OpenStudy (ash2326):

\[5^2=x\] We know that \[log_x y=z=> y^z=x\] \[5^2=x=> \log_ 5 x=2\]

OpenStudy (anonymous):

Equivalent of \(5^2=x\) logs give the exponent, so 2 is the exponent. 5 is the base. Thus; \[\log_5 x=2\] For \(\log_4 64\), write 64 as a power of 4. \[\log_4 4^3\] Once again, since logs give exponents, and the bases are equal, take the exponent as your answer, 3.

OpenStudy (ash2326):

Sorry \[log_x y=z=> x^z=y\]

OpenStudy (anonymous):

thankyou everyone !!!!!!!!!!